# DDAD: A Two-pronged Adversarial Defense Based on Distributional Discrepancy

### v1

@@ E.7 Combination with Adversarial Training
E.8 Compute Resources
Table 11: Training time (hours : minutes : seconds) and memory consumption (MB) for DDAD on CIFAR-10 and ImageNet-1K . This table reports the compute resources for the entire training process of DDAD (i.e., optimizing MMD + training the denoiser).
Dataset GPU Batch Size Classifier Training Time Memory
CIFAR-10 2 $\times$ NVIDIA A100 500 RN-18 00:28:17 5927
WRN-28-10 00:55:34 6276
ImageNet-1K 4 $\times$ NVIDIA A100 128 RN-50 09:52:50 97354
Table 12: Inference time (hours : minutes : seconds) for DDAD on CIFAR-10 and ImageNet-1K. This table reports the comput resources for evaluating the entire test set of CIFAR-10 (i.e., 10,000 images) and ImageNet-1K (i.e., 50,000 images).
Dataset GPU Batch Size Classifier Inference Time
CIFAR-10 1 $\times$ NVIDIA A100 100 WRN-28-10 00:00:32
ImageNet-1K 2 $\times$ NVIDIA A100 100 RN-50 00:03:08
Table 11 presents the compute resources for DDAD, which include GPU configurations, batch size, classifier, training time, and memory usage for each dataset. For CIFAR-10, using 2 NVIDIA A100 GPUs with a batch size of 500, our method’s training time is approximately 28 minutes with ResNet-18 and 55 minutes with WideResNet-28-10. The memory consumption is 5927 MB and 6276 MB, respectively. For ImageNet-1K, using 4 NVIDIA A100 GPUs with a batch size of 128, our method’s training time is approximately 10 hours, with a memory consumption of 97354 MB. Compared to AT baseline methods, DDAD offers better training efficiency (e.g., it can scale to large datasets like ImageNet-1K). This is mainly because we directly use the pre-trained classifier. Furthermore, training MMD is extremely fast (usually less than 1 minute on CIFAR-10) and we use a lightweight denoiser.
Table 12 presents the compute resources for evaluating DDAD, which include GPU configurations, batch size, classifier and inference time for each dataset. For CIFAR-10, using 1 NVIDIA A100 GPU with a batch size of 100, our method’s inference time is approximately 32 seconds over the entire test set of CIFAR-10. For ImageNet-1K, using 2 NVIDIA A100 GPUs with a batch size of 100, our method’s inference time is approximately 3 minutes over the entire test set of ImageNet-1K. Although DDAD requires training an extra denoiser and MMD-OPT, it significantly outperforms AP baselines in inference speed. Furthermore, relying on a pre-trained generative model is not always feasible, as training such models at scale can be highly resource-intensive. Therefore, considering considering the trade-off between computational cost and the performance of DDAD, we believe that training an additional detector and denoiser is feasible and worthwhile. In general, DDAD provides a more lightweight design.
Appendix F Detailed Related Work
@@
Adversarial attacks. The discovery of adversarial examples (AEs) has raised a security concern for AI development in recent decades (Szegedy et al., 2014; Goodfellow et al., 2015). AEs are often crafted by adding imperceptible noise to clean images, which can easily mislead a classifier to make wrong predictions. The algorithms that generate AEs are called adversarial attacks. For example, the Fast Gradient Sign Method (FGSM) involves adding noise to the clean data in the direction of the gradient of the loss function with respect to the clean data (Goodfellow et al., 2015). Expanding on FGSM, the Basic Iterative Method (BIM) (Kurakin et al., 2017) iteratively applies small noises to the clean data in the direction of the gradient of the loss function, updating the input at each step to create more effective AEs than single-step methods such as FGSM. Madry et al. (2018) propose the Projected Gradient Descent (PGD), which further improves the iterative approach of BIM by adding random initialization to the input data before applying iterative gradient-based perturbations. Beyond non-targeted attacks, the Carlini & Wagner attack (C&W) specifically directs data towards a chosen target label, which crafts AEs by optimizing a specially designed objective function (Carlini & Wagner, 2017). AutoAttack (AA) (Croce & Hein, 2020a) is an ensemble of multiple adversarial attacks, which combines three non-target white-box attacks (Croce & Hein, 2020b) and one targeted black-box attack (Andriushchenko et al., 2020), which makes AA a benchmark standard for evaluating adversarial robustness. However, the computational complexity of AA is relatively high. Gao et al. (2022) propose the Minimum-margin attack (MMA), which can be used as a faster alternative to AA. Beyond computing exact gradients, Athalye et al. (2018b) propose Expectation over Transformation (EOT) to correctly compute the gradient for defenses that apply randomized transformations to the input. Athalye et al. (2018a) propose the Backward Pass Differentiable Approximation (BPDA), which approximates the gradient with an identity mapping to effectively break the defenses that leverage obfuscated gradients. According to Lee & Kim (2023), PGD+EOT is currently the best attack for denoiser-based defense methods.
@@
Adversarial training. Another prominent defensive framework is adversarial training (AT). Vanilla AT (Madry et al., 2018) directly generates and incorporates AEs during the training process, forcing the model to learn the underlying distributions of AEs. Besides vanilla AT, several modifications have been developed to enhance the effectiveness of AT. For instance, at the early stage of AT, Song et al. (2019) propose to treat adversarial attacks as a domain adaptation problem and enhance the generalization of AT by minimizing the distributional discrepancy. Zhang et al. (2019) propose optimizing a surrogate loss function based on theoretical bounds. Similarly, Wang et al. (2020) explore how misclassified examples influence a model’s robustness, leading to an improved adversarial risk through regularization. From the perspective of reweighting, Ding et al. (2020) propose to reweight adversarial data with instance-dependent perturbation bounds $\epsilon$ and Zhang et al. (2021) introduce a geometry-aware instance-reweighted AT (GAIRAT) framework, which differentiates weights based on the proximity of data points to the class boundary. Wang et al. (2021) builds upon GAIRAT by leveraging probabilistic margins to reweight AEs due to their continuous nature and independence from specific perturbation paths. More recently, Zhang et al. (2024) propose to pixel-wisely reweight AEs by explicitly guiding them to focus on important pixel regions. Other modifications include improving AT using data augmentation methods (Gowal et al., 2021; Rebuffi et al., 2021) and hyper-parameter selection methods (Gowal et al., 2020; Pang et al., 2021). Although AT achieves high robustness against particular attacks, it suffers from significant degradation in clean accuracy and high computational complexity (Wong et al., 2020; Laidlaw et al., 2021; Poursaeed et al., 2021). Different from the AT framework, our method does not train a robust classifier. Instead, by directly feeding detected CEs to a pre-trained classifier, our method can effectively maintain clean accuracy. Meanwhile, by using a lightweight detector and denoiser model, our method can alleviate the computational complexity.
@@
Denoiser-based adversarial defense. Another well-known defense framework is denoiser-based adversarial defense, which often leverages generative models to shift AEs back to their clean counterparts before feeding them into a classifier. In most literature, it is called adversarial purification (AP). Previous methods mainly focus on exploring the use of more powerful generative models for AP. For example, at the early stage of AP, Meng & Chen (2017) propose a two-step process called MagNet to remove adversarial noise by first using a detector to discard the detected AEs, and then leveraging the reconstructability of an autoencoder to purify the rest of the examples, which guides AEs towards the manifold of clean data. After MagNet, Liao et al. (2018) design a denoising UNet that can denoise AEs to their clean counterparts by reducing the distance between adversarial and clean data under high-level representations. Samangouei et al. (2018) use a GAN trained on clean examples to project AEs onto the generator’s manifold. Song et al. (2018) find that AEs lie in low-probability regions of the image distribution and propose to maximize the probability of a given test example. Naseer et al. (2020) focus on training a conditional GAN, which engages in a min-max game with a critic network, to differentiate between adversarial and clean data. Yoon et al. (2021) propose to use the denoising score-based model to purify adversarial examples. Nie et al. (2022) propose to use diffusion models to remove adversarial noise by gradually adding Gaussian noise to AEs, and then wash out the noise by solving the reverse-time stochastic differential equation. The success of recent AP methods often relies on the assumption that there will be a pre-trained generative model that can precisely estimate the probability density of the CEs (Yoon et al., 2021; Nie et al., 2022). However, even powerful generative models (e.g., diffusion models) may have an inaccurate density estimation, leading to unsatisfactory performance (Chen et al., 2024). By contrast, instead of estimating probability densities, our method directly minimizes the distributional discrepancies between AEs and CEs, leveraging the fact that identifying distributional discrepancies is simpler and more feasible than estimating density. Nayak et al. (2023) propose to use MMD as a regularizer during the training of the denoiser. Different from their work, we use an optimized version of MMD (i.e., MMD-OPT), which is more sensitive to adversarial attacks. Furthermore, the MMD-OPT serves not only as a ‘guider’ during training to help minimize the distributional discrepancy between AEs and CEs, but also a ‘detector’ that helps distinguish AEs and CEs.
Appendix G Discussions on Batch-wise Detections
@@
Benefits of using batch-wise statistical hypothesis test. A main benefit of using a batch-wise statistical hypothesis test is that it can effectively control the false alarm rate. For example, for DDAD, we set the maximum false alarm rate to be 5%. Fang et al. (2022) theoretically prove that for instance-wise detection methods to work perfectly, there must be a gap in the support set between IID and out-of-distribution (OOD) data. This theory also applies to adversarial problems, but such a support set does not exist in adversarial settings, making perfect instance-wise detection generally difficult.
@@
Limitation and its solutions for user inference. DDAD leverages statistics based on distributional discrepancies (i.e., MMD-OPT), which requires the data to be processed in batches for adversarial detection. However, when the batch size is too small, the stability of DDAD will be affected (see Figure 2). To address this issue, for user inference, single samples provided by the user can be dynamically stored in a queue. Once the queue accumulates enough samples to form a batch, our method can then process the batch collectively using the proposed approach. A direct cost of this solution is the waiting time, as the system must accumulate enough samples (e.g., 50 samples) to form a batch before processing. However, in scenarios where data arrives quickly, the waiting time is typically very short, making this approach feasible for many real-time applications. For applications with stricter latency requirements, the batch size can be dynamically adjusted based on the incoming data rate to minimize waiting time. For instance, if the system detects a lower data arrival rate, it can process smaller batches to ensure timely responses. Overall, it is a trade-off problem: using our method for user inference can obtain high robustness, but the cost is to wait for batch processing. Based on the performance improvements our method obtains over the baseline methods, we believe the cost is feasible and acceptable. Another possible solution is to find more robust statistics that can measure distributional discrepancies with fewer samples. Recently, measuring the expected score of a sample after multiple perturbations has proven useful for this purpose (Zhang et al., 2023). However, computing the expected score is time-consuming. We emphasize that this paper primarily focuses on the relationship between distributional discrepancies and adversarial risk, aiming to inspire the design of a new defense method. We leave it as future work.

### v2

@@ D.8 Combination with Adversarial Training
Appendix E Detailed Related Work
@@
Adversarial attacks. The discovery of adversarial examples (AEs) has raised a security concern for AI development in recent decades (Szegedy et al., 2014; Goodfellow et al., 2015). AEs are often crafted by adding imperceptible noise to clean images, which can easily mislead a classifier to make wrong predictions. The algorithms that generate AEs are called adversarial attacks. For example, the Fast Gradient Sign Method (FGSM) perturbs clean data in the direction of the loss gradient (Goodfellow et al., 2015). Expanding on FGSM, the Basic Iterative Method (BIM) (Kurakin et al., 2017) iteratively applies small noises to the clean data in the direction of the gradient of the loss function, updating the input at each step to create more effective AEs than single-step methods such as FGSM. Madry et al. (2018) propose the Projected Gradient Descent (PGD), which further improves the iterative approach of BIM by adding random initialization to the input data before applying iterative gradient-based perturbations. Beyond non-targeted attacks, the Carlini & Wagner attack (C&W) specifically directs data towards a chosen target label, which crafts AEs by optimizing a specially designed objective function (Carlini & Wagner, 2017). AutoAttack (AA) (Croce & Hein, 2020a) is an ensemble of multiple adversarial attacks, which combines three non-target white-box attacks (Croce & Hein, 2020b) and one targeted black-box attack (Andriushchenko et al., 2020), which makes AA a benchmark standard for evaluating adversarial robustness. However, the computational complexity of AA is relatively high. Gao et al. (2022) propose the Minimum-margin attack (MMA), which can be used as a faster alternative to AA. Beyond computing exact gradients, Athalye et al. (2018b) propose Expectation over Transformation (EOT) to correctly compute the gradient for defenses that apply randomized transformations to the input. Athalye et al. (2018a) propose the Backward Pass Differentiable Approximation (BPDA), which approximates the gradient with an identity mapping to effectively break the defenses that leverage obfuscated gradients. According to Lee & Kim (2023), PGD+EOT is currently the best attack for denoiser-based defense methods.
@@
Adversarial training. Another prominent defensive framework is adversarial training (AT). Vanilla AT (Madry et al., 2018) directly generates and incorporates AEs during the training process, forcing the model to learn the underlying distributions of AEs. Besides vanilla AT, several modifications have been developed to enhance the effectiveness of AT. For instance, at the early stage of AT, Song et al. (2019) propose to treat adversarial attacks as a domain adaptation problem and enhance the generalization of AT by minimizing the distributional discrepancy. Zhang et al. (2019) propose optimizing a surrogate loss function based on theoretical bounds. Similarly, Wang et al. (2020) explore how misclassified examples influence a model’s robustness, leading to an improved adversarial risk through regularization. From the perspective of reweighting, Ding et al. (2020) propose to reweight adversarial data with instance-dependent perturbation bounds $\epsilon$ and Zhang et al. (2021) introduce a geometry-aware instance-reweighted AT (GAIRAT) framework, which differentiates weights based on the proximity of data points to the class boundary. Wang et al. (2021) build upon GAIRAT by leveraging probabilistic margins to reweight AEs due to their continuous nature and independence from specific perturbation paths. Zhou et al. (2023b) propose a joint adversarial defense method that combines a phase-level adversarial training mechanism to enhance robustness against phase-based attacks with an amplitude-based preprocessing operation to mitigate perturbations in the amplitude domain. More recently, Zhang et al. (2024) propose to pixel-wisely reweight AEs by explicitly guiding them to focus on important pixel regions. Other modifications include improving AT using data augmentation methods (Gowal et al., 2021; Rebuffi et al., 2021) and hyper-parameter selection methods (Gowal et al., 2020; Pang et al., 2021). Although AT achieves high robustness against particular attacks, it suffers from significant degradation in clean accuracy and high computational complexity (Wong et al., 2020; Laidlaw et al., 2021; Poursaeed et al., 2021). Different from the AT framework, our method does not train a robust classifier. Instead, by directly feeding detected CEs to a pre-trained classifier, our method can effectively maintain clean accuracy. By using a lightweight detector and denoiser model, our method can alleviate the computational complexity.
@@
Denoiser-based adversarial defense. Another well-known defense framework is denoiser-based adversarial defense, which often leverages generative models to shift AEs back to their clean counterparts before feeding them into a classifier. In most literature, it is called adversarial purification (AP). Previous methods mainly focus on exploring the use of more powerful generative models for AP. For example, at the early stage of AP, Meng & Chen (2017) propose a two-step process called MagNet, which first discards detected AEs using a detector, then uses an autoencoder to purify the rest by guiding them toward the manifold of clean data. After MagNet, Liao et al. (2018) design a denoising UNet that can denoise AEs to their clean counterparts by reducing the distance between adversarial and clean data under high-level representations. Samangouei et al. (2018) use a GAN trained on clean examples to project AEs onto the generator’s manifold. Song et al. (2018) find that AEs lie in low-probability regions of the image distribution and propose to maximize the probability of a given test example. Naseer et al. (2020) focus on training a conditional GAN, which engages in a min-max game with a critic network, to differentiate between adversarial and clean data. Yoon et al. (2021) propose to use the denoising score-based model to purify adversarial examples. Nie et al. (2022) propose to use diffusion models to remove adversarial noise by gradually adding Gaussian noise to AEs, and then wash out the noise by solving the reverse-time stochastic differential equation. Zhou et al. (2023a) propose to utilize complementary masks to disrupt adversarial noise and employs guided denoising models to recover robust and predictive representations from the masked samples. The success of recent AP methods often relies on the assumption that there will be a pre-trained generative model that can precisely estimate the probability density of the CEs (Yoon et al., 2021; Nie et al., 2022). However, even powerful generative models (e.g., diffusion models) may have an inaccurate density estimation, leading to unsatisfactory performance (Chen et al., 2024). By contrast, instead of estimating probability densities, our method directly minimizes the distributional discrepancies between AEs and CEs, leveraging the fact that identifying distributional discrepancies is simpler and more feasible than estimating density. Nayak et al. (2023) propose to use MMD as a regularizer during the training of the denoiser. Different from their work, we use an optimized version of MMD (i.e., MMD-OPT), which is more sensitive to adversarial attacks. Furthermore, the MMD-OPT not only acts as a guiding signal to minimize the distributional discrepancy between AEs and CEs, but also acts as a discriminator to distinguish between CEs and AEs, which kills two birds with one stone.
Appendix F Discussions on Batch-wise Detections
@@
Benefits of using batch-wise statistical hypothesis test. A main benefit of using a batch-wise statistical hypothesis test is that it can effectively control the false alarm rate. For example, for DAD, we set the maximum false alarm rate to be 5%. Fang et al. (2022) theoretically prove that for instance-wise detection methods to work perfectly, there must be a gap in the support set between IID and out-of-distribution (OOD) data. This theory also applies to adversarial problems, but such a support set does not exist in adversarial settings, making perfect instance-wise detection generally difficult.
@@
Limitation and its solutions for user inference. DAD leverages statistics based on distributional discrepancies (i.e., MMD-OPT), which requires the data to be processed in batches for adversarial detection. However, when the batch size is too small, the stability of DAD will be affected (see Figure 2). To address this issue, for user inference, single samples provided by the user can be dynamically stored in a queue. Once the queue accumulates enough samples to form a batch, our method can then process the batch collectively using the proposed approach. A direct cost of this solution is the waiting time, as the system must accumulate enough samples (e.g., 50 samples) to form a batch before processing. However, in scenarios where data arrives quickly, the waiting time is typically very short, making this approach feasible for many real-time applications. For applications with stricter latency requirements, the batch size can be dynamically adjusted based on the incoming data rate to minimize waiting time. For instance, if the system detects a lower data arrival rate, it can process smaller batches to ensure timely responses. Overall, it is a trade-off problem: using our method for user inference can obtain high robustness, but the cost is to wait for batch processing. Based on the performance improvements our method obtains over the baseline methods, we believe the cost is feasible and acceptable. Another possible solution is to find more robust statistics that can measure distributional discrepancies with fewer samples. Recently, measuring the expected score of a sample after multiple perturbations has proven useful for this purpose (Zhang et al., 2023). However, computing the expected score is time-consuming. We emphasize that this paper primarily focuses on the relationship between distributional discrepancies and adversarial risk, aiming to inspire the design of a new defense method. We leave it as future work.